OpenStudy (anonymous):

Solving Linear Equations 3x-(x-5)=-2x

OpenStudy (er.mohd.amir):

open braket which change sign and trasfer x one side and number other side

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

i dont understand i thought i multiply the outside of the side to the inside of the ( )

OpenStudy (er.mohd.amir):

3x-x+5=-2x

OpenStudy (skullpatrol):

– ( x – 5 ) = (-1) • ( x – 5 )

OpenStudy (anonymous):

ok i got it but can you answer how to i know when to open the brackets because sometimes it have to be muti from the outside to the inside of the ( )

OpenStudy (skullpatrol):

-x + 5

OpenStudy (skullpatrol):

If there is a minus sign outside of the brackets you have to convert it into a -1 and use the distributive property :-)

OpenStudy (jiteshmeghwal9):

First step is to open the brackets 3x-x+5=-2x now collect the like terms one side 2x+2x =-5 4x=-5

OpenStudy (anonymous):

-5/4

OpenStudy (skullpatrol):

a(b+c) = ab + ac

OpenStudy (jiteshmeghwal9):

yes x=-5/4.

OpenStudy (skullpatrol):

Now substitute -5/4 back into the original equation to check if you got it right :-)